OpenStudy (anonymous):

Find the product of (x + 7)^2. Is the answer x^2 + 49 ? @jim_thompson5910

OpenStudy (anonymous):

@eSpeX @radar

jimthompson5910 (jim_thompson5910):

no, that's a very common misconception

jimthompson5910 (jim_thompson5910):

(x + 7)^2 = (x+7)(x+7) distribute or use FOIL to expand

OpenStudy (anonymous):

I forgot how to do the Foil thingy > . <

jimthompson5910 (jim_thompson5910):

FOIL = first, outer, inner, last

jimthompson5910 (jim_thompson5910):

FOIL F: multiply first terms: x*x = x^2 O: multiply outer terms: x*7 = 7x I: multiply inner terms: 7*x = 7x L: multiply last terms: 7*7 = 49 then you add/combine like terms x^2 + 14x + 49

OpenStudy (anonymous):

Oh, Thanks ^_^

jimthompson5910 (jim_thompson5910):

yw